OpenStudy (anonymous):

If the length of a large banner is 4x + 5 and the width is half as long, what is the area? A. 6x² + 20x + 12.5 B. 8x² + 17x + 12.5 C. 8x² + 20x + 12.5 D. 8x² + 22x + 12.5

OpenStudy (anonymous):

do 4x+5 times 4x+5/2

OpenStudy (anonymous):

4x+5 times (4x+5)/2 i mean

OpenStudy (studygurl14):

\[(4x + 5) \times [\frac{ (4x + 5) }{ 2 }] = A\]

OpenStudy (anonymous):

Area is =length*width given len=(4x+5)*{(4x+5)/2}=.......,